Calculate the area of a triangle whose vertices are given by (3,-1,2),(1,-1,2),(4,-2,1)
1
Expert's answer
2019-03-15T12:45:45-0400

Let


"A=(3,-1,2),\quad B=(1,-1,2), \quad C=(4,-2,1)"

So


"\overrightarrow{AB}=(1-3,-1-(-1),2-2)=(-2,0,0)\\
\overrightarrow{AC}=(4-3,-2-(-1),1-2)=(1,-1,-1)"


The vector product


"[\overrightarrow{AB}\times \overrightarrow{AC}]=\begin{vmatrix}
 \hat i & \hat j & \hat k\\
 -2 & 0 & 0\\
1 & -1 & -1
\end{vmatrix}=-2\hat j+2\hat k"

The area of the triangle

"A=\frac{1}{2}|[\overrightarrow{AB}\times \overrightarrow{AC}]|=\frac{1}{2}\sqrt{(-2)^2+2^2}=\sqrt{2}"